# Does the Hamiltonian determine the tensor product structure and the 3d space?

### 04 Jan 2024 v2

@@ Abstract
These results affect some proposals to recover the 3d space from the Hamiltonian.
@@ II Proof of non-uniqueness
Any TPS obtained uniquely only from the Hamiltonian’s spectrum should be preserved by any unitary symmetry transformation $\widehat{\mathbf{S}}$ preserving the Hamiltonian, $\widehat{\mathbf{S}}\widehat{\mathbf{H}}=\widehat{\mathbf{H}}\widehat{\mathbf{S}}$. Otherwise, the method to obtain the TPS is based on additional assumptions, introducing more information than the Hamiltonian’s spectrum. To see this, consider that the method to obtain the TPS consists of calculations done in a basis of $\mathcal{H}$ in which $\widehat{\mathbf{H}}$ is diagonal, with the eigenvalues sorted in ascending order. If the resulting TPS depends only on the spectrum of $\widehat{\mathbf{H}}$, we should obtain the same TPS if we use a different basis in which $\widehat{\mathbf{H}}$ has the same form. Any such basis can be obtained from the former basis by a unitary transformation that commutes with $\widehat{\mathbf{H}}$.
The TPS can be specified uniquely in terms of algebras of observables $\mathcal{A}_{1},\mathcal{A}_{2},\ldots$, i.e. algebras of operators on $\mathcal{H}$, required to generate the full algebra of observables of $\mathcal{H}$, and to satisfy the conditions [11]
$\left[\mathcal{A}_{j},\mathcal{A}_{k}\right]=\widehat{\mathbf{0}}\textnormal{ and }\mathcal{A}_{j}\cap\mathcal{A}_{k}=\mathbb{C}\widehat{\mathbf{1}}_{\mathcal{H}}\textnormal{ for all }j\neq k.$ (2)
@@
$\big{(}\mathcal{A}^{\prime}_{j}\big{)}_{j\in\{1,\ldots,n\}}=\big{(}\widehat{\mathbf{S}}\mathcal{A}_{j}\widehat{\mathbf{S}}^{\dagger}\big{)}_{j\in\{1,\ldots,n\}}.$ (3)
If the TPS were unique, this would imply that, for any $\widehat{\mathbf{S}}$ satisfying $\widehat{\mathbf{S}}\widehat{\mathbf{H}}=\widehat{\mathbf{H}}\widehat{\mathbf{S}}$, the two sets of algebras of observables are identical,
$\left\{\mathcal{A}^{\prime}_{j};j=1,\ldots,n\right\}=\left\{\widehat{\mathbf{S}}\mathcal{A}_{j}\widehat{\mathbf{S}}^{\dagger};j=1,\ldots,n\right\}.$ (4)
@@
The TPS is invariant under unitary transformations of $\mathcal{H}$ acting restrictively on each of the factor Hilbert spaces $\mathcal{H}_{j}$, and under the permutations of equal dimension factors $\mathcal{H}_{j}$ [11, 3]. The group of local unitary transformations acting only on $\mathcal{H}_{j}$ is $\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}$. But since $\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}\cap\textnormal{U}\big{(}\mathcal{H}_{k}\big{)}=\textnormal{U}(1)$ for $j\neq k$, local phase changes are equivalent to the global ones. Taking this (and the fact that permutations of the factors don’t affect the dimension) into account, it follows that
$\dim\textnormal{U}_{\textnormal{TPS}}$ $\leq\sum_{j=1}^{n}\big{(}\dim\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}-\underbrace{\dim\textnormal{U}(1)}_{\textnormal{local}}\big{)}+\underbrace{\dim\textnormal{U}(1)}_{\textnormal{global}}\big{)}$ (6)
@@
If the necessary condition (5) were satisfied, any maximal commutative subgroup of $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ would be included in a maximal commutative subgroup of $\textnormal{U}_{\textnormal{TPS}}$. I will show that this doesn’t happen.
The dimension of any maximal commutative subgroup of $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ is at most
@@
The dimension of any maximal commutative subgroup of $\textnormal{U}_{\textnormal{TPS}}$ is at most
@@
$d_{1}\cdot\ldots\cdot d_{n}>\sum_{j=1}^{n}d_{j}-n+1.$ (10)
Then, the necessary condition (5) can’t be satisfied, and there are symmetries preserving the Hamiltonian that don’t preserve the TPS. Therefore, the TPS can’t be uniquely determined by the Hamiltonian’s spectrum.
@@
Let’s see how fast does the number of additional parameters needed grows as $n$ grows. If all subsystems are qubits or qudits of the same dimension $d_{j}=d$, the dimensions of the two groups $\textnormal{U}_{\textnormal{TPS}}$ and $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ are
$\begin{array}[]{lll}\dim&\textnormal{U}_{\textnormal{TPS}}&=1+n(d^{2}-1),\\ \dim&\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}&\geq d^{n}.\\ \end{array}$ (11)
@@ III Discussion: TPS, entanglement, and emergent spacetime
Now we’ll see that, even if the math would have allowed some Hamiltonians to determine a unique TPS, such a Hamiltonian would not generate a unitary dynamics able to entangle subsystems. Of course, we now know from Theorem 1 that the TPS can’t be unique, but it’s interesting to explore some physical reasons why this idea couldn’t work to begin with. This will also help seeing why we can’t declare the resulting TPSs to be equivalent in any physically meaningful way.
The operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$, $t\in\mathbb{R}$, commute with $\widehat{\mathbf{H}}$, so if the TPS is unique, when they are used as symmetry transformations, they should preserve it. But they also act as evolution operators of the system,
@@
One may still hope that the TPS, while not unique, it is unique up to non-physical symmetry transformations. For example that non-uniqueness can be removed by taking equivalence classes of TPSs. But then again, the entanglement properties of such a system would not be able to change.
One may be distracted by using the operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ to prove that a unique TPS prevents entanglement, and propose that the TPS is determined up to the one-parameter group of symmetries $\big{(}e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}\big{)}_{t\in\mathbb{R}}$. This may even seem like a nice feature, because then the method would seem to recover not something about the system at a fixed $t$, but something about its entire history. But this would be a red herring: I only chose the operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ as an example, a useful one since we know that the entanglement can be changed by unitary time evolution. But in fact there are many other one-parameter groups that commute with $\widehat{\mathbf{H}}$, each of them being able to generate TPSs with different entanglement invariants, and therefore they can’t be equivalent.
It was proposed that, if a TPS emerges uniquely from the spectrum, the factors of the TPS can be identified with the points of the space, or whatever discrete building blocks of space we think should replace points in a quantum theory of gravity [2, 1].
The idea to recover the continuous $3$d-space from discrete building blocks is clever and it makes sense [5], and indeed these building blocks correspond to tensor product structures.
Then, if the TPS were unique, the distances could be recovered by using information that is not contained in the spectrum of the Hamiltonian, but that can be extracted from the state $\widehat{\mathbf{\rho}}$ and its relation with the TPS. Then maybe the $3$d space manifold structure would emerge at large scales. Carroll and Singh use the mutual information for this [2, 1].
@@
But more importantly, if the building blocks of space correspond to the factors of the TPS, Theorem 1 implies that they can’t emerge uniquely from the spectrum. Additional observables (along with their physical meaning [9]) are needed to begin with.

### 08 Jan 2024 v3

@@ Abstract
These results affect the proposals to recover the 3d space from the Hamiltonian.
@@ II Proof of non-uniqueness
Any TPS obtained uniquely only from the Hamiltonian’s spectrum should be preserved by any unitary symmetry transformation $\widehat{\mathbf{S}}$ preserving the Hamiltonian, $\widehat{\mathbf{S}}\widehat{\mathbf{H}}\widehat{\mathbf{S}}^{\dagger}=\widehat{\mathbf{H}}$. Otherwise, this indicates that the method to obtain the TPS is based on additional hidden assumptions, introducing more information than the Hamiltonian’s spectrum. To see this, consider that the method to obtain the TPS consists of calculations done in a basis of $\mathcal{H}$ in which $\widehat{\mathbf{H}}$ is diagonal, with the eigenvalues sorted in ascending order. If the resulting TPS depends only on the spectrum of $\widehat{\mathbf{H}}$, we should obtain the same TPS if we use a different basis in which $\widehat{\mathbf{H}}$ has the same form. Any such basis can be obtained from the former basis by a unitary transformation that commutes with $\widehat{\mathbf{H}}$.
The TPS can be specified uniquely in terms of algebras of observables $\mathcal{A}_{1},\mathcal{A}_{2},\ldots$, i.e. algebras of operators on $\mathcal{H}$, required to generate the full algebra of observables of $\mathcal{H}$, and to satisfy, for all $j\neq k$, the conditions [13]
$\left[\mathcal{A}_{j},\mathcal{A}_{k}\right]=\widehat{\mathbf{0}}\textnormal{ and }\mathcal{A}_{j}\cap\mathcal{A}_{k}=\mathbb{C}\widehat{\mathbf{1}}_{\mathcal{H}}.$ (2)
@@
$\left\{\mathcal{A}^{\prime}_{1},\mathcal{A}^{\prime}_{2},\ldots\right\}=\left\{\widehat{\mathbf{S}}\mathcal{A}_{1}\widehat{\mathbf{S}}^{\dagger},\widehat{\mathbf{S}}\mathcal{A}_{2}\widehat{\mathbf{S}}^{\dagger},\ldots\right\}.$ (3)
If the TPS were unique, this would imply that, for any $\widehat{\mathbf{S}}$ satisfying $\widehat{\mathbf{S}}\widehat{\mathbf{H}}\widehat{\mathbf{S}}^{\dagger}=\widehat{\mathbf{H}}$, the two sets of algebras of observables are identical,
$\left\{\mathcal{A}_{1},\mathcal{A}_{2},\ldots\right\}=\left\{\widehat{\mathbf{S}}\mathcal{A}_{1}\widehat{\mathbf{S}}^{\dagger},\widehat{\mathbf{S}}\mathcal{A}_{2}\widehat{\mathbf{S}}^{\dagger},\ldots\right\}.$ (4)
@@
The TPS is invariant under unitary transformations of $\mathcal{H}$ acting restrictively on each of the factor Hilbert spaces $\mathcal{H}_{j}$, and under the permutations of the factors $\mathcal{H}_{j}$ of the same dimension [13, 3]. The group of local unitary transformations acting only on $\mathcal{H}_{j}$ is $\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}$. But since $\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}\cap\textnormal{U}\big{(}\mathcal{H}_{k}\big{)}=\textnormal{U}(1)$ for $j\neq k$, local phase changes are equivalent to the global ones. Taking this (and the fact that group of permutations of the factors doesn’t affect the dimension) into account, it follows that
$\dim\textnormal{U}_{\textnormal{TPS}}$ $\leq\underbrace{\dim\textnormal{U}(1)}_{\textnormal{global}}+\sum_{j=1}^{n}\big{(}\dim\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}-\underbrace{\dim\textnormal{U}(1)}_{\textnormal{local}}\big{)}$ (6)
@@
If the necessary condition (5) were satisfied, any commutative subgroup of $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ would be included in a commutative subgroup of $\textnormal{U}_{\textnormal{TPS}}$. I will show that this doesn’t happen.
Since $\dim\mathcal{H}=d_{1}\cdot\ldots\cdot d_{n}$, the dimension of any commutative subgroup of $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ is at most
@@
But the dimension of any commutative subgroup of $\textnormal{U}_{\textnormal{TPS}}$ is at most
@@
$d_{1}\cdot\ldots\cdot d_{n}>\sum_{j=1}^{n}d_{j}-n+1,$ (10)
so $\textnormal{D}_{\widehat{\mathbf{H}}}>\textnormal{D}_{\textnormal{TPS}}.$ Then, the necessary condition (5) can’t be satisfied, and there are symmetries preserving the Hamiltonian that don’t preserve the TPS. Therefore, the TPS can’t be uniquely determined by the Hamiltonian’s spectrum.
@@
Let’s see how fast the number of additional parameters needed grows as $n$ grows. If all subsystems are qubits or qudits of the same dimension $d_{j}=d$, the dimensions of the two groups $\textnormal{U}_{\textnormal{TPS}}$ and $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ are
$\begin{array}[]{lll}\dim&\textnormal{U}_{\textnormal{TPS}}&=n(d^{2}-1)+1,\\ \dim&\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}&\geq d^{n}.\\ \end{array}$ (11)
@@ III Discussion: TPS, entanglement, and emergent spacetime
Now we’ll see that, even if the math would have allowed some Hamiltonians to determine a unique TPS, such a Hamiltonian would not generate a unitary dynamics able to entangle subsystems. Of course, we now know from Theorem 1 that the TPS can’t be unique, but it’s interesting to explore some physical reasons why this idea couldn’t work to begin with. This will also help see that, even if we would hope to get rid of non-uniqueness by taking equivalence classes of the resulting TPSs, this would not make sense physically.
The operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$, $t\in\mathbb{R}$, commute with $\widehat{\mathbf{H}}$, so if the TPS were unique, when they are used as symmetry transformations, they should preserve it. But they also act as evolution operators of the system,
@@
I chose the operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ as an example, a useful one since we know that the entanglement can be changed by unitary time evolution. But there are infinitely many other one-parameter groups that commute with $\widehat{\mathbf{H}}$, each of them being able to generate TPSs with different entanglement invariants, and therefore they can’t be equivalent.
One may think to avoid the existence of the numerous non-equivalent TPSs implied by the fact that entanglement properties can change, by declaring all of them equivalent. That is, declare that all TPSs obtained from one another by the symmetry transformations $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ are physically the same TPS. But the entropy of entanglement is relative to the TPS, and any such equivalence class of TPSs would contain TPSs with respect to which the entanglement properties are different. So any TPS constructed by such an equivalence would be unable to exhibit changes in the entropy of entanglement, because all these changes would stay within the same equivalence class. Different states with different entanglement properties would be declared to be physically equivalent. And we know that in reality the entropy of entanglement can be changed by unitary evolution, and different states can have different entanglement properties. So whatever we get from a construction of the TPS from the spectrum and an equivalence relation, it would not be the TPS that we know from physics.
It was proposed that, if a TPS emerges uniquely from the spectrum, the factors of the TPS can be identified with the points of the space, or whatever discrete building blocks of space we think should replace points in a quantum theory of gravity [2, 1]. The idea to recover the continuous $3$d-space from discrete building blocks is clever and it makes sense [5], and indeed these building blocks correspond to tensor product structures.
Then, if the TPS were unique, the distances between the resulting points of space could be recovered by using information that is not contained in the spectrum of the Hamiltonian, but that can be extracted from the state $\widehat{\mathbf{\rho}}$ and its relation with the TPS. Then maybe the $3$d space manifold structure would emerge at large scales. Carroll and Singh use the mutual information for this [2, 1].
@@
But more importantly, if the building blocks of space correspond to the factors of the TPS, Theorem 1 implies that they can’t emerge uniquely from the spectrum. For the infinite-dimensional Hilbert space case, infinitely many examples of systems with the same Hamiltonian and unit vector, but with different TPSs and different dimensions of space are already known to exist [9]. Moreover, even if the TPS is given along with the Hamiltonian and a unit vector, this is insufficient to describe our world [11]. Additional observables (along with their physical meaning [10]) are needed to begin with.